OpenStudy (goformit100):

The total number of digits used in numbering the pages of a book having 366 pages is A. 732 B. 990 C. 1098 D. 1305

OpenStudy (goformit100):

@Noemi95

OpenStudy (anonymous):

366 pages...

OpenStudy (anonymous):

the first 9 pages use 1 digit each, the next how many pages use 2 digits each, and the next how many pages use 3 digits each?

OpenStudy (goformit100):

Madam @Noemi95

OpenStudy (anonymous):

is this question specifically for nemo?

OpenStudy (anonymous):

B.990 1 - 9 = 1*9 digits 10 -99 = 2*10*9 = 180 digits 100-366 = 3* 267 = 801 digits total: 990

OpenStudy (anonymous):

@Noemi95 please don't give the answer right away...

OpenStudy (anonymous):

I just did! (>.\)

OpenStudy (anonymous):

@Peter14

OpenStudy (anonymous):

I know. please don't. explain how to find the answer.

OpenStudy (goformit100):

I AM SORRY @Peter14

OpenStudy (anonymous):

You are right @Peter14, I usually do that. @goformit100 I think you get the overall concept of all your questions by now (if there's more) Try to work it out on your own, and if you can't, just let me know and I'll walk you through the steps. (: